Consider an imaginary economy that has been growing at a rate of 4% per year. Government economists have proposed a number of policies to increase the growth rate but first need to convince the President that the policies will pay off. To do so, they want to present a comparison of the number of years it will take for the economy to double, depending on the growth rate.

According to the rule of 70, determine the number of years it will take the economy to double at each growth rate.

Growth Rate Years Required to Double
(Percent) (Nearest whole number of years)
4
5
6
Answer #1

According to the rule of 70, the approximate number of years n for a variable growing at the constant growth rate of R percent, to double is equal to 70 divided by rate of groeth r .

n= 70/r

for r= 4% doubling time, n = 70/4 = 17.5

for r= 5% doubling time, n = 70/5 = 14

for r= 6% doubling time, n = 70/6 = 11.666

Growth Rate Years Required to Double
(Percent) (Nearest whole number of years)
4 17.5
5 14
6 11.66666667